# Balancing LoRA Performance and Efficiency with Simple Shard Sharing

### v9

@@ Abstract
Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA), effectively reduce the number of trainable parameters in Large Language Models (LLMs). However, as model scales continue to grow, the demand for computational resources remains a significant challenge. Existing LoRA variants often struggle to strike an optimal balance between adaptability (model performance and convergence speed) and efficiency (computational overhead, memory usage, and initialization time). This paper introduces Fossil(Framework for Optimal Shard Sharing Integration in LoRA), a novel PEFT approach that addresses this trade-off through a simple shard-sharing mechanism. Fossil leverages the insight that a low-rank adaptation can be achieved by decomposing the weight matrix into multiple fragment matrices and utilizing a shared, trainable common fragment. This method constructs the low-rank update matrix through the replication of these shared, partitioned shards. We also propose a hardware-efficient and broadly applicable implementation for Fossil. Extensive experiments conducted on a range of tasks, alongside a systematic analysis of computational performance, demonstrate Fossil’s superiority. The results show that Fossil significantly outperforms standard LoRA and its prominent variants in both model performance metrics and computational efficiency, including initialization speed and training throughput. By effectively balancing expressive power and resource utilization, Fossil offers a compelling solution for efficiently adapting large-scale models.
@@ 1 Introduction
Fine-tuning Large Language Models (LLMs) [16, 17] is a prevalent methodology for adapting these models to specific downstream tasks. However, the comprehensive fine-tuning of all parameters typically incurs substantial computational costs. Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques [23] have been developed to mitigate the training expenditure associated with these large-scale models. Among such techniques, Low-Rank Adaptation (LoRA) [8] has distinguished itself as one of the most prominent PEFT methods. LoRA employs a low-rank approximation for the weight updates, a strategy that offers a markedly reduced number of tunable parameters, notable efficacy when compared to full fine-tuning, and the potential for zero inference overhead. LoRA constructs this low-rank adaptation matrix through an intuitive design, positing that the weight update $\Delta\bm{W}$ can be approximated by the product of two lower-rank matrices, $\bm{B}\bm{A}\approx\Delta\bm{W}$. Evidently, this specific factorization is not necessarily the optimal low-rank approximation of the original $\Delta\bm{W}$.
@@
Considering the aforementioned trade-off, our investigation reveals that a surprisingly simple strategy—constructing a low-rank matrix through the replication of a shared, partitioned shard—can achieve a compelling balance across these criteria. This approach, which we introduce as Fossil,  exhibits notable strengths. In terms of adaptability, Fossil demonstrates superior approximation fidelity compared to LoRA and several of its variants. Its relatively substantial gradient norm (as illustrated in Figure 4) promotes rapid convergence, a key characteristic of its high adaptability. Furthermore, concerning efficiency, Fossil benefits from rapid initialization; its hardware-friendly, repetition-based initialization facilitates swift processing, particularly when scaling to large models with substantial parameter counts. The formulation of Fossil can also be expressed in an efficient and mathematically equivalent form that exhibits favorable time and space complexity.
@@
Fossil(Ours) $y=\bm{W_{0}}x+\operatorname{expand}(\bm{D})x$ $\bm{D}\sim 0$
Fossile(Ours) $y=\bm{W_{0}}x+\bm{D^{\top}S}$ $\bm{D}\sim 0$
We perform scaling experiments to verify our preliminary conclusions. Our extensive evaluation shows that Fossil demonstrates excellent performance on understanding and generation tasks, significantly surpassing other baseline method. Moreover, we present the training dynamics i.e., loss curves, which Fossil also showcase a significantly faster convergence rate compared to other baselines. In generation tasks, Fossil demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variants across the board. This demonstrates the feasibility of the framework. Furthermore, we validate the efficiency of Fossil  which enables it to significantly outperform LoRA and its variants in both computational efficiency and memory usage.
@@
    In Section 4, we proposed Fossil, an efficient and adaptable structure with shard sharing mechanism.
@@
Figure 1: Left. Schematic diagram of the forward computation and initialization methods in LoRA and Fossil. Right. PyTorch-style pseudocode illustrating the implementation of Fossil.
@@ 3 No Free Launch: Balancing Between Adaptability and Efficiency
We considered diverse methods  [^1] We have not included methods such as LoRA-GA [21] or LoRA+ [7] in our current analysis. While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider Fossil to be largely orthogonal to them. Consequently, the analytical techniques employed in their study may still offer valuable insights for Fossil.: (1) Full-parameter finetuning [13]. (2) LoRA [8]. (3) Alternatives to LoRA w/ different architectures, including: PiSSA [14], VeRA [11], DoRA [12] and MoRA [10]. (4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA [19], MoS [22]. (1) An overview of their forward form, initialization method can be found at Table 1.
@@ Results.
Figure 2 illustrates the comparative adaptability of different methods. We utilize the minimum validation loss achieved by each approach as an indicator of its expressive capacity when approximating the performance of full-parameter fine-tuning. The results clearly demonstrate that methods leveraging singular value decomposition (SVD), such as PiSSA, attain a relatively low loss. Conversely, efficiency-focused techniques like MoS exhibit higher losses. A plausible explanation for this discrepancy is that such methods further decompose LoRA matrices into shared components which may inherently constrain their expressive power. Our method Fossil reaches a relatively advanced performance comparing to other variants.
@@
The efficacy (See Figure 2) of Fossil is evident: its strategic combination of parameter sharing and an efficient computational design culminates in rapid, scalable performance across both initialization and training stages. In contrast, while techniques like PiSSA demonstrate commendable adaptability, as shown in prior experiments, their reliance on computationally intensive Singular Value Decomposition for initialization significantly hampers their overall speed. Other approaches, such as VeRA and AdaLoRA, offer efficient initialization and computation; however, as previously discussed, they often achieve this at the cost of comparatively reduced adaptability.
4 Fossil: Shard Sharing for the Performance and Efficiency Tradeoff
@@ 4.1 Method Overview
Based on this insight, we propose Fossil, which defines the weight update $\Delta\bm{W}$ as a large matrix generated from a small trainable matrix $\bm{D}$ through an expansion operation. The updating of $\bm{W}$ and the forward pass can be expressed as:
@@
4.2 Efficient Implementation of Fossil
The above formulation is effective in the initialization process, as it only needs to initialize a small $\bm{D}$. However, directly computing $\operatorname{expand}(\bm{D})x$ has a time complexity of $O(bldk)$ and memory complexity of $O(dk)$ , which can be computationally intensive. It is obvious that Fossil can be transformed into an efficient form that leverages the block structure of the input to avoid explicitly forming the large matrix, by redefining $\bm{D}\in\mathbb{R}^{r\times d}$, where $r$ is a tunable rank parameter. Instead of partitioning the output dimension $d$, we divide the input dimension $k$ into $r$ blocks, each of size $g=\lfloor k/r\rfloor$ (for simplicity, assume $k$ is divisible by $r$). For an input $\bm{x}\in\mathbb{R}^{b\times l\times k}$, partition it along the $k$-dimension, and sum each block along the $k$-dimension:
@@ 4.2 Efficient Implementation of Fossil
The efficient implementation of Fossil relies on an innovative input aggregation mechanism, namely blockwise input summation. We highlight its advantages through the following steps: (1) Input Partitioning and Aggregation: The aggregation exploits local redundancy in the input, preserving critical information while reducing the computational dimensionality. (2) Fast Computation: The cost of computing the efficient form is significantly lower than the original complexity. (3) Resource Savings: Memory usage drops comparing to original form. For example, with $k=1024$ and $r=16$, memory is reduced by about 64 times. An overall analysis of space and time complexity is analyzed in Table 5.
@@
Figure 3: Left. Fossil consumes less GPU memory and achieves higher throughput(# tokens per second). GPU memory usage refers to the amount of memory occupied by an RTX 4090 card. Right. The results of fine-tuning LLaMA2-7B on MetaMathQA.
@@ 5 Experiments
In this section, we evaluate the performance of Fossil on various benchmark datasets. Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset with the robert-base model. Subsequently, we evaluated the Natural Language Generation (NLG [^2] Evaluations were performed with the OpenCompass repository, and the Math dataset was evaluated using a 5-shot prompt configuration.) capabilities by fine-tuning the LLM. To ensure fair comparisons, we select methods officially included in the PEFT repository to avoid discrepancies caused by inconsistent tuning implementations.
Table 2: The results of fine-tuning RoBERTa-base using Fossil and various LoRA variants were compared on a subset of the GLUE benchmark.
@@
Fossil 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60
Table 3: We conducted fine-tuning of large language models using Fossil and multiple LoRA variants, evaluating their performance on GSM8K, Math, HumanEval, and MBPP benchmarks. All reported results are averaged over three independent runs to ensure robustness. The first-place entry should be highlighted in bold, and the second-place entry should be underlined.
@@ 5.1 Experiments on Natural Language Understanding
We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC. Performance is evaluated on the development set using accuracy as the primary metric. The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU. Each experiment is conducted with 3 different random seeds, and the average performance is reported. As shown in Table 2, Fossil demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and PiSSA.
@@ 5.2 Experiment on Natural Language Generation
To verify the generalizability of Fossil, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math and code. (1) Math: We trained our model on a 395k subset of MetaMathQA [25], a dataset bootstrapped from other math instruction tuning datasets like GSM8K [5] and MATH  [25], with higher complexity and diversity. (2) Code: We train our model on a 100k subset of CodeFeedback [26], a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval [4] and Mbpp [1]. The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used. It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing Fossil with LoRA. To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of Fossil to 64, ensuring that Fossil uses fewer parameters than LoRA to demonstrate its superiority. Each experiment is conducted with 2 different random seeds, and the average performance is reported.
32641282560612182430rankgradient normftLoRAPiSSAFossil
Figure 4: Left. The loss curve for LLaMA2-7B fine-tuned on MetaMathQA shows Fossil’s superior fitting ability across architectures and parameter settings, with a notably rapid loss reduction in the first 100 steps, underscoring its effectiveness. Right. The initial gradient norm of model training.
@@ Results.
As shown in Table 6, Fossil consistently achieves state-of-the-art performance across multiple models and evaluation metrics. Specifically, on LLaMA2-7B, Fossil surpasses PiSSA on most metrics while using fewer trainable parameters than both LoRA and DoRA, showcasing its parameter efficiency and superior performance. On RWKV6-7B, where all methods share the same number of trainable parameters, Fossil still achieves the best overall performance, indicating its strong adaptation capability to non-transformer architectures. Across all evaluated models, including Mistral-7B, LLaMA2-13B, and Qwen3-4B, Fossil maintains its leading position on key benchmarks, often achieving the highest scores with significantly fewer parameters—further highlighting its effectiveness and scalability.
@@ 5.3 Effect of Rank $r$
This subsection explores the upper limits of the Fossil structure by varying the rank $r$ in the Fossil matrix. Comparative experiments were conducted by fine-tuning LLaMA2-7B  [^3] We use LLaMA2-7B instead of LLaMA3-8B because we found that the LLaMA3 series is over-optimized on math-related tasks. on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when $r=16$, the Fossil structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is only 3.73. The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of $r$ impacts the model’s ability to understand unseen data. Based on this observation, we hypothesize that when the rank is too small, it significantly limits the model’s generalization ability.
Table 4: Comparing different values of rank $(r)$ on LLaMA2-7B with Fossil.
@@
Fossil $O(dr)$ $O(bldk)$
Fossile $O(dr)$ $O(blr(d+\frac{k}{r}))$

### v10

@@ Abstract
Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA), effectively reduce the number of trainable parameters in Large Language Models (LLMs). However, as model scales continue to grow, the demand for computational resources remains a significant challenge. Existing LoRA variants often struggle to strike an optimal balance between adaptability (model performance and convergence speed) and efficiency (computational overhead, memory usage, and initialization time). This paper introduces MiSS(Matrix Shard Sharing ), a novel PEFT approach that addresses this trade-off through a simple shard-sharing mechanism. MiSS leverages the insight that a low-rank adaptation can be achieved by decomposing the weight matrix into multiple fragment matrices and utilizing a shared, trainable common fragment. This method constructs the low-rank update matrix through the replication of these shared, partitioned shards. We also propose a hardware-efficient and broadly applicable implementation for MiSS. Extensive experiments conducted on a range of tasks, alongside a systematic analysis of computational performance, demonstrate MiSS’s superiority. The results show that MiSS significantly outperforms standard LoRA and its prominent variants in both model performance metrics and computational efficiency, including initialization speed and training throughput. By effectively balancing expressive power and resource utilization, MiSS offers a compelling solution for efficiently adapting large-scale models.
@@ 1 Introduction
Fine-tuning Large Language Models (LLMs) [16, 17, 25] is a prevalent methodology for adapting these models to specific downstream tasks. However, the comprehensive fine-tuning of all parameters typically incurs substantial computational costs. Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques [23] have been developed to mitigate the training expenditure associated with these large-scale models. Among such techniques, Low-Rank Adaptation (LoRA) [8] has distinguished itself as one of the most prominent PEFT methods. LoRA employs a low-rank approximation for the weight updates, a strategy that offers a markedly reduced number of tunable parameters, notable efficacy when compared to full fine-tuning, and the potential for zero inference overhead. LoRA constructs this low-rank adaptation matrix through an intuitive design, positing that the weight update $\Delta\bm{W}$ can be approximated by the product of two lower-rank matrices, $\bm{B}\bm{A}\approx\Delta\bm{W}$. Evidently, this specific factorization is not necessarily the optimal low-rank approximation of the original $\Delta\bm{W}$.
@@
Considering the aforementioned trade-off, our investigation reveals that a surprisingly simple strategy—constructing a low-rank matrix through the replication of a shared, partitioned shard—can achieve a compelling balance across these criteria. This approach, which we introduce as MiSS,  exhibits notable strengths. In terms of adaptability, MiSS demonstrates superior approximation fidelity compared to LoRA and several of its variants. Its relatively substantial gradient norm (as illustrated in Figure 4) promotes rapid convergence, a key characteristic of its high adaptability. Furthermore, concerning efficiency, MiSS benefits from rapid initialization; its hardware-friendly, repetition-based initialization facilitates swift processing, particularly when scaling to large models with substantial parameter counts. The formulation of MiSS can also be expressed in an efficient and mathematically equivalent form that exhibits favorable time and space complexity.
@@
MiSS(Ours) $y=\bm{W_{0}}x+\operatorname{expand}(\bm{D})x$ $\bm{D}\sim 0$
MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D^{\top}S}$ $\bm{D}\sim 0$
We perform scaling experiments to verify our preliminary conclusions. Our extensive evaluation shows that MiSS demonstrates excellent performance on understanding and generation tasks, significantly surpassing other baseline method. Moreover, we present the training dynamics i.e., loss curves, which MiSS also showcase a significantly faster convergence rate compared to other baselines. In generation tasks, MiSS demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variants across the board. This demonstrates the feasibility of the framework. Furthermore, we validate the efficiency of MiSS  which enables it to significantly outperform LoRA and its variants in both computational efficiency and memory usage.
@@
    In Section 4, we proposed MiSS, an efficient and adaptable structure with shard sharing mechanism.
@@
Figure 1: Left. Schematic diagram of the forward computation and initialization methods in LoRA and MiSS. Right. PyTorch-style pseudocode illustrating the implementation of MiSS.
@@ 3 No Free Launch: Balancing Between Adaptability and Efficiency
We considered diverse methods  [^1] We have not included methods such as LoRA-GA [21] or LoRA+ [7] in our current analysis. While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider MiSS to be largely orthogonal to them. Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning [13]. (2) LoRA [8]. (3) Alternatives to LoRA w/ different architectures, including: PiSSA [14], VeRA [11], DoRA [12] and MoRA [10]. (4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA [19], MoS [22]. (1) An overview of their forward form, initialization method can be found at Table 1.
@@ Results.
Figure 2 illustrates the comparative adaptability of different methods. We utilize the minimum validation loss achieved by each approach as an indicator of its expressive capacity when approximating the performance of full-parameter fine-tuning. The results clearly demonstrate that methods leveraging singular value decomposition (SVD), such as PiSSA, attain a relatively low loss. Conversely, efficiency-focused techniques like MoS exhibit higher losses. A plausible explanation for this discrepancy is that such methods further decompose LoRA matrices into shared components which may inherently constrain their expressive power. Our method MiSS reaches a relatively advanced performance comparing to other variants.
@@
The efficacy (See Figure 2) of MiSS is evident: its strategic combination of parameter sharing and an efficient computational design culminates in rapid, scalable performance across both initialization and training stages. In contrast, while techniques like PiSSA demonstrate commendable adaptability, as shown in prior experiments, their reliance on computationally intensive Singular Value Decomposition for initialization significantly hampers their overall speed. Other approaches, such as VeRA and AdaLoRA, offer efficient initialization and computation; however, as previously discussed, they often achieve this at the cost of comparatively reduced adaptability.
4 MiSS: Shard Sharing for the Performance and Efficiency Tradeoff
@@ 4.1 Method Overview
Based on this insight, we propose MiSS, which defines the weight update $\Delta\bm{W}$ as a large matrix generated from a small trainable matrix $\bm{D}$ through an expansion operation. The updating of $\bm{W}$ and the forward pass can be expressed as:
@@
4.2 Efficient Implementation of MiSS
The above formulation is effective in the initialization process, as it only needs to initialize a small $\bm{D}$. However, directly computing $\operatorname{expand}(\bm{D})x$ has a time complexity of $O(bldk)$ and memory complexity of $O(dk)$ , which can be computationally intensive. It is obvious that MiSS can be transformed into an efficient form that leverages the block structure of the input to avoid explicitly forming the large matrix, by redefining $\bm{D}\in\mathbb{R}^{r\times d}$, where $r$ is a tunable rank parameter. Instead of partitioning the output dimension $d$, we divide the input dimension $k$ into $r$ blocks, each of size $g=\lfloor k/r\rfloor$ (for simplicity, assume $k$ is divisible by $r$). For an input $\bm{x}\in\mathbb{R}^{b\times l\times k}$, partition it along the $k$-dimension, and sum each block along the $k$-dimension:
@@ 4.2 Efficient Implementation of MiSS
The efficient implementation of MiSS relies on an innovative input aggregation mechanism, namely blockwise input summation. We highlight its advantages through the following steps: (1) Input Partitioning and Aggregation: The aggregation exploits local redundancy in the input, preserving critical information while reducing the computational dimensionality. (2) Fast Computation: The cost of computing the efficient form is significantly lower than the original complexity. (3) Resource Savings: Memory usage drops comparing to original form. For example, with $k=1024$ and $r=16$, memory is reduced by about 64 times. An overall analysis of space and time complexity is analyzed in Table 5.
@@
Figure 3: Left. MiSS consumes less GPU memory and achieves higher throughput(# tokens per second). GPU memory usage refers to the amount of memory occupied by an RTX 4090 card. Right. The results of fine-tuning LLaMA2-7B on MetaMathQA.
@@ 5 Experiments
In this section, we evaluate the performance of MiSS on various benchmark datasets. Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset with the robert-base model. Subsequently, we evaluated the Natural Language Generation (NLG [^2] Evaluations were performed with the OpenCompass repository, and the Math dataset was evaluated using a 5-shot prompt configuration.) capabilities by fine-tuning the LLM. To ensure fair comparisons, we select methods officially included in the PEFT repository to avoid discrepancies caused by inconsistent tuning implementations.
Table 2: The results of fine-tuning RoBERTa-base using MiSS and various LoRA variants were compared on a subset of the GLUE benchmark.
@@
MiSS 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60
Table 3: We conducted fine-tuning of large language models using MiSS and multiple LoRA variants, evaluating their performance on GSM8K, Math, HumanEval, and MBPP benchmarks. All reported results are averaged over three independent runs to ensure robustness. The first-place entry should be highlighted in bold, and the second-place entry should be underlined.
@@ 5.1 Experiments on Natural Language Understanding
We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC. Performance is evaluated on the development set using accuracy as the primary metric. The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU. Each experiment is conducted with 3 different random seeds, and the average performance is reported. As shown in Table 2, MiSS demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and PiSSA.
@@ 5.2 Experiment on Natural Language Generation
To verify the generalizability of MiSS, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math and code. (1) Math: We trained our model on a 395k subset of MetaMathQA [26], a dataset bootstrapped from other math instruction tuning datasets like GSM8K [5] and MATH  [26], with higher complexity and diversity. (2) Code: We train our model on a 100k subset of CodeFeedback [27], a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval [4] and Mbpp [1]. The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used. It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing MiSS with LoRA. To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of MiSS to 64, ensuring that MiSS uses fewer parameters than LoRA to demonstrate its superiority. Each experiment is conducted with 2 different random seeds, and the average performance is reported.
32641282560612182430rankgradient normftLoRAPiSSAMiSS
Figure 4: Left. The loss curve for LLaMA2-7B fine-tuned on MetaMathQA shows MiSS’s superior fitting ability across architectures and parameter settings, with a notably rapid loss reduction in the first 100 steps, underscoring its effectiveness. Right. The initial gradient norm of model training.
@@ Results.
As shown in Table 6, MiSS consistently achieves state-of-the-art performance across multiple models and evaluation metrics. Specifically, on LLaMA2-7B, MiSS surpasses PiSSA on most metrics while using fewer trainable parameters than both LoRA and DoRA, showcasing its parameter efficiency and superior performance. On RWKV6-7B, where all methods share the same number of trainable parameters, MiSS still achieves the best overall performance, indicating its strong adaptation capability to non-transformer architectures. Across all evaluated models, including Mistral-7B, LLaMA2-13B, and Qwen3-4B, MiSS maintains its leading position on key benchmarks, often achieving the highest scores with significantly fewer parameters—further highlighting its effectiveness and scalability.
@@ 5.3 Effect of Rank $r$
This subsection explores the upper limits of the MiSS structure by varying the rank $r$ in the MiSS matrix. Comparative experiments were conducted by fine-tuning LLaMA2-7B  [^3] We use LLaMA2-7B instead of LLaMA3-8B because we found that the LLaMA3 series is over-optimized on math-related tasks. on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when $r=16$, the MiSS structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is only 3.73. The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of $r$ impacts the model’s ability to understand unseen data. Based on this observation, we hypothesize that when the rank is too small, it significantly limits the model’s generalization ability.
Table 4: Comparing different values of rank $(r)$ on LLaMA2-7B with MiSS.
@@
MiSS $O(dr)$ $O(bldk)$
MiSSe $O(dr)$ $O(blr(d+\frac{k}{r}))$